# Aegis: A Decentralized Expansion Blockchain

### v2

@@ 4.2 Chain Structure
All nodes start with a single Genesis Aegis block (Algorithm 4.3 line 3). All Aegis blocks $b$ (apart from Genesis) have a parent Aegis block $\textit{parent}(b)$ and an Aegis-to-Primary reference to a primary block $\textit{ref}_{A\to P}(b)$. (For genesis only: $\textit{parent}(b_{\textit{genesis}})=\bot$ and $\textit{ref}_{A\to P}(b_{\textit{genesis}})=\bot$.)
@@
3      if A checkpoint in $\underline{\mathcal{L}}$ conflicts with $b$ then return False
@@ 14                $)\mathcal{N}\leftarrow\textit{stakers}(\textit{ref}_{A\to P}(b)$
1 on Receive at time $t$ entry $e$ with pointed-to block $b$ and its parent $b^{\prime}$
@@
           assert no entries later than $t-\Delta_{\textit{active}}$
@@
                     $B_{\textit{reset}}\leftarrow\underline{\mathcal{L}}(\textit{ref}_{A\to P}^{\>\textit{reset}}(b)$)
                      (Primary-chain reset entry)
@@ 14                          $B_{\textit{reset}}\leftarrow\bot$
4.3 Primary-Chain Contract
The primary-chain contract (Algorithm 2) allows checkpointing Aegis blocks and initiating resets. It receives entries submitted by Aegis nodes asking to register either a checkpoint or a reset.
It registers a reset if there are no entries in the last $\Delta_{\textit{active}}$ steps, indicating one is indeed necessary; no other checks are required.
@@ 3 $\mathcal{B}_{i}\leftarrow\mathcal{B}_{i}\cup\textup{blocks received}$
4 ¯$\underline{\mathcal{L}}_{i}\leftarrow\underline{\mathcal{L}}_{i}\|\textup{Primary chain extension receieved}$ (No conflicts by assumption) $B_{\textit{last}}\leftarrow$ last contract entry in ¯$\underline{\mathcal{L}}_{i}$ or $\bot$ if none
5 if $B_{\textit{last}}$ is a reset  then
@@
7           $B_{\textit{checkpoint}}\leftarrow$ latest checkpoint, $\bot$ if none
@@ 9           $t_{0}\leftarrow t_{\textit{gen}}(B_{\textit{last}})$
16                     else if $B_{\textit{last}}$ is a checkpoint  then
@@ 17                          $B_{\textit{checkpoint}}\leftarrow B_{\textit{last}}$
                          $\mathcal{N}\leftarrow\textit{stakers}(\underline{\mathcal{L}}(\textit{ref}_{A\to P}(\mathcal{B}_{i}(\textit{ref}_{P\to A}(B_{\textit{checkpoint}})))))$
@@
                          $t_{0}\leftarrow t_{\textit{gen}}(\underline{\mathcal{L}}(\textit{ref}_{A\to P}(\mathcal{B}_{i}(\textit{ref}_{P\to A}(B_{\textit{checkpoint}})))))$
@@ 44           $\mathcal{N}\leftarrow\textit{stakers}(\textit{ref}_{A\to P}(b))$
48           if $t<t_{0}+\Delta_{\textit{active}}-3\Delta_{\underline{\mathcal{L}}\textit{-write}}$ then (Latest committee has enough time)
49                if $b=b_{\textit{checkpointed}}\land B_{\textit{last}}$ is reset then (No blocks since last checkpoint and following a reset)
50                     $b\leftarrow\textit{consensusStep}(\textup{propose a block},\bigl{(}\mathcal{L}(|\mathcal{L}|-1),B_{\textit{last}}\bigr{)},\mathcal{N})$
51
@@
53                          $b\leftarrow\textit{consensusStep}(\textup{propose a block},\bigl{(}\mathcal{L}(|\mathcal{L}|-1),\bot\bigr{)},\mathcal{N})$
54
@@
62
                                     (No-op if $t_{0}+\Delta_{\textit{active}}-\Delta_{\underline{\mathcal{L}}\textit{-write}}<t<t_{0}+\Delta_{\textit{active}}$)
@@ 4.4 Node Protocol
4.4.1 Sync to Latest Checkpoint
At the beginning of a step a node first reads the states of the primary and expansion blockchains (lines 33). Then, it finds the latest checkpointed block as follows. If the latest primary-chain entry is a checkpoint, it takes this checkpoint; the committee for the next block to log is expected to be the one defined by the checkpointed block, i.e., according to the primary block it references (line 3). Otherwise, if the latest entry is a reset and the committee defined by the reset is still active (line 3), it takes the most recent checkpoint with the reset’s committee for the subsequent block. If the reset’s committee is inactive, the node issues a new reset and returns, waiting for it to take place. Finally, if there is no primary block entry then this is the first execution of the protocol, so the node issues a reset and returns, waiting for it to take place.
@@ 4.4.1 Sync to Latest Checkpoint
4.4.2 Sync Past Checkpoint
@@ 4.4.3 Extend the log
4.5 Forensics
Aegis takes advantage of the consensus protocol’s forensics support [54, 20] to penalize Byzantine nodes if their number is greater than its threshold and they successfully violate agreement. The forensics procedure encompasses both the contract and the peer-to-peer protocol, we describe it only here for pseudocode readability.
@@ 4.5 Forensics
5 Security Overview
The proof is deferred to Appendix B; we briefly review it here. To prove security we must show that Aegis nodes do not register different values at the same height (Agreement), that if all have the same value and none is Byzantine they decide that value (Validity), that after $t_{\textit{GST}}$ they register values frequently (Progress), and that violation leads to penalties.
We first show by induction that the ancestors of a block generated by an active committee were also generated by active committees. Then, using an indistinguishability argument we show that the guarantees of an existing consensus protocol (with an infinitely-running committee) are preserved during the validity period of an Aegis committee.
The proofs of the three properties follow a similar structure. We prove by induction on the Aegis block number. For each we show the property holds either due to the consensus guarantees, or due to the interaction of the nodes protocol and the primary-chain contract.
6 Conclusion
We present Aegis, a novel expansion blockchain protocol whose nodes can join and leave in a permissionless fashion by staking in a primary chain. Aegis guarantees safety at all times and progress after $t_{\textit{GST}}$ despite asynchronous communication among its nodes (up to $t_{\textit{GST}}$), using committees dynamically specified by a primary-chain smart contract with synchronous access.
While Aegis can be deployed as-is on top of a primary blockchain, there are two practical aspects to consider. Aegis requires primary-chain transactions in intervals close to the time for unstaking $\Delta_{\textit{active}}$. This implies a tradeoff between short unstaking time and low primary-chain overhead. In terms of performance, Aegis can change committee membership at a lower frequency than every block. This allows using contemporary BFT state machine replication algorithms (e.g., [64, 28, 7]) that are more efficient than running a one-shot consensus instance per block.
The tools to deploy staking for Aegis are ready at hand, e.g., Avalanche’s elastic subnets [38]; direct implementation with a smart contract; and Ethereum’s EigenLayer [60], which enables recycling of large stake amounts. Thus, Aegis opens the door for a new generation of truly decentralized expansion blockchains.
Appendix A Notation
Table 1 summarizes key notation used in this paper.
Table 1: Notation
Symbol Meaning
$\alpha$ Minimum ratio of correct nodes in an active committee
$t_{\textit{GST}}$ Global Stabilization Time
$\Delta_{\textit{prop}}$ Network propagation bound after $t_{\textit{GST}}$
$H(B)=\hat{B}$ Hash of block $B$
$\underline{\mathcal{L}}$ Primary ledger
$\mathcal{L}$ Aegis ledger
$\mathcal{L}(k)$ Aegis block at height $k$
$\textit{parent}(b)$ Parent block of block $b$
$\mathcal{B}(\hat{b})$ Block $b$ if $b$ is in block set $B$, otherwise $\bot$
$\textit{ref}_{A\to P}(b)$ Reference from block $b$ to a primary block
$\textit{stakers}(B)$ Committee defined by primary block $B$
$t_{\textit{gen}}(b)$ Generation time of block $b$
Appendix B Security
After showing that the ancestors of a block generated by an active committee were also generated by active committees (§B.1), we prove Aegis achieves agreement (§B.2), validity (§B.3) and progress (§B.4), and that it penalizes misbehaving nodes (§B.5). The security guarantees follow (§B.6).
B.1 Active Committee Follows Active Committees
We call a committee active if its nodes’ collateral was locked recently enough so it could not be withdrawn yet.
@@ B.1 Active Committee Follows Active Committees
B.2 Agreement
For agreement we need two more lemmas. First, we ensure steps taken by the nodes are reflected in the primary chain. We show that if the nodes agree on the ledger up to some height, then a subsequent block will be accepted by the primary-chain contract.
@@ B.2 Agreement
B.3 Validity
To prove validity, we first note that the consensus protocol guarantees validity of an active committee decision.
@@ B.3 Validity
The proof is similar to that of Lemma 8. We are now ready to prove validity.
Proposition 11 (Validity).
Aegis achieves validity.
@@
Assumption. For $k>0$, assume that the claim holds for ${k-1}$. That is, for all nodes $i$ and times $t$, if node $i$ logs a block $b$ at height $k^{\prime}\leq k-1$ at time $t_{1}$ ($\mathcal{L}_{i}^{t_{1}}(k^{\prime})\neq\bot$) and all nodes are correct with input $v_{k^{\prime}}$ for height $k^{\prime}$, then ${\mathcal{L}_{i}^{t_{1}}(k^{\prime})=v_{k^{\prime}}}$. Also assume there is no checkpoint for a block $b$ at a height $k^{\prime}\leq k-1$ with value different from $v_{k^{\prime}}$.
Step. Now we prove for $k$. Assume for contradiction that there exists a nodes $i$ and time $t$ such that node $i$ logs a block $b$ in position $k$ at time $t$ and $\mathcal{L}_{i}^{t_{1}}(k)=b\neq v_{k}$.
@@
In both cases, that committee decided on an invalid value, contradicting the consensus protocol agreement property (Lemma 10).
@@ B.4 Progress
If at time $t$ the Aegis checkpoint is later than $t-\Delta_{\textit{active}}+\Delta_{\underline{\mathcal{L}}\textit{-write}}$, the nodes take consensus steps and decide within $\Delta_{\textit{consensus}}$ steps (Algorithm 4.3 line 4, Lemma 12). However, if the consensus runs for too long, and its result might not be checkpointed in time, it is abandoned (Algorithm 4.3 line 4). Then no progress can be made until $\Delta_{\textit{active}}$ after the last checkpoint/reset and a new committee is instated with a new reset (line 4). The reset is registered within $\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps and a new committee reaches consensus within $\Delta_{\textit{consensus}}$ (Lemma 12), guaranteed after $t_{\textit{GST}}$ because ${\Delta_{\textit{consensus}}<\Delta_{\textit{active}}-\Delta_{\underline{\mathcal{L}}\textit{-write}}}$.
@@
Note. Independent Proof-of-Stake chains are vulnerable to long-range attacks, where a deprecated committee forms an alternative history long after it unstaked. Joining nodes cannot distinguish between real history and the forged one. Aegis is not vulnerable to this attack due to the security of the primary chain.
B.5 Slashing
Enforcing the desired behavior of correct nodes while they are active necessitates penalizing them if they successfully violate safety (Agreement or Validity). We show Aegis achieves this.
@@ B.5 Slashing
If a node $i$ belongs to an active committee that violates the consensus safety, it will be penalized.
@@
B.6 Aegis Security
The conclusion follows directly from propositions 911, and 13, since nodes are penalized (Lemma 14) for violating safety.
@@ B.6 Aegis Security
Aegis ensures Agreement, validity, and progress.

### v3

@@ 5.2. Chain Structure
All nodes start with a single Genesis Aegis block. All Aegis blocks $b$ (apart from Genesis) have a parent Aegis block $\textit{parent}(b)$ and an Aegis-to-Primary reference to a primary block $\textit{ref}_{A\to P}(b)$. (For genesis only: $\textit{parent}(b_{\textit{genesis}})=\bot$ and $\textit{ref}_{A\to P}(b_{\textit{genesis}})=\bot$.)
@@ 13                $)\mathcal{N}\leftarrow\textit{stakers}(\textit{ref}_{A\to P}(b)$
1 on Receive at time $t$ entry $e$ referencing Aegis block $b$ and its parent $b^{\prime}$
@@
           assert no entries in blocks later than $t-\Delta_{\textit{active}}$
@@
                     $B_{\textit{reset}}\leftarrow$ primary block containing reset entry at $\textit{ref}_{A\to P}^{\>\textit{reset}}(b)$
                      (Primary-chain block with reset entry)
@@ 14                          $B_{\textit{reset}}\leftarrow\bot$
Both Aegis blocks and primary blocks have associated generation times. For any block $b$, we denote its generation time by $t_{\textit{gen}}(b)$, representing the timestamp when the block was created. In the protocol, only the timestamps of primary blocks are used to determine committee activity periods and to enforce timing constraints to avoid potential vulnerabilities via timestamp manipulation of Aegis blocks.
5.3. Primary-Chain Contract
The primary-chain contract (Algorithm 2) allows checkpointing Aegis blocks and initiating resets. It receives entries submitted by Aegis nodes asking to register either a checkpoint or a reset. In each block the primary-chain contract may receive one entry at most, and ignores any further submissions.
The contract registers a reset if there are no entries in the last $\Delta_{\textit{active}}$ steps, indicating one is indeed necessary; no other checks are required.
@@ 3 $\mathcal{B}_{i}\leftarrow\mathcal{B}_{i}\cup\textup{blocks received}$
4 ¯$\underline{\mathcal{L}}_{i}\leftarrow\underline{\mathcal{L}}_{i}\|\textup{Primary chain extension receieved}$ (No conflicts by assumption) $B_{\textit{last}}\leftarrow$ primary block containing the last contract entry in ¯$\underline{\mathcal{L}}_{i}$ or $\bot$ if none
5 if entry in $B_{\textit{last}}$ is a reset  then
@@
7           $B_{\textit{checkpoint}}\leftarrow$ primary block containing the latest checkpoint entry, $\bot$ if none
@@ 9           $t_{0}\leftarrow t_{\textit{gen}}(B_{\textit{last}})$
16                     else if entry in $B_{\textit{last}}$ is a checkpoint  then
@@ 17                          $B_{\textit{checkpoint}}\leftarrow B_{\textit{last}}$
18                          $b_{\textit{checkpointed}}\leftarrow\mathcal{B}_{i}(\textit{ref}_{P\to A}(B_{\textit{checkpoint}}))$
                           (The checkpointed Aegis block)
                          $\mathcal{N}\leftarrow\textit{stakers}(\underline{\mathcal{L}}(\textit{ref}_{A\to P}(b_{\textit{checkpointed}})))$
@@
                          $t_{0}\leftarrow t_{\textit{gen}}(\underline{\mathcal{L}}(\textit{ref}_{A\to P}(b_{\textit{checkpointed}})))$
@@ 45           $\mathcal{N}\leftarrow\textit{stakers}(\textit{ref}_{A\to P}(b))$
48
49           if $t<t_{0}+\Delta_{\textit{active}}-3\Delta_{\underline{\mathcal{L}}\textit{-write}}$  then (Latest committee has enough time to extend)
50                if $b=b_{\textit{checkpointed}}\land$ entry in $B_{\textit{last}}$ is a reset then (No blocks since last checkpoint and following a reset)
51                     $b\leftarrow\textit{consensusStep}(v^{i}_{t}(|\mathcal{L}|),\bigl{(}\mathcal{L}(|\mathcal{L}|-1),B_{\textit{last}}\bigr{)},\mathcal{N})$
52
@@
54                          $b\leftarrow\textit{consensusStep}(v^{i}_{t}(|\mathcal{L}|),\bigl{(}\mathcal{L}(|\mathcal{L}|-1),\bot\bigr{)},\mathcal{N})$
55
@@
63
                                     (No-op if $t_{0}+\Delta_{\textit{active}}-3\Delta_{\underline{\mathcal{L}}\textit{-write}}\leq t\leq t_{0}+\Delta_{\textit{active}}$)
@@ 5.4. Node Protocol
5.4.1. Sync to Latest Checkpoint
At the beginning of a step a node first reads the states of the primary and tethered blockchains (lines 33). Then, it finds the latest checkpointed block as follows. If the latest primary-chain entry is a checkpoint, it takes this checkpoint; the committee for the next block to log is expected to be the one defined by the checkpointed block, i.e., according to the primary block it references (line 3). Otherwise, if the latest entry is a reset and the committee defined by the reset is still active (line 3), it takes the most recent checkpoint with the reset’s committee for the subsequent block. If the reset’s committee is inactive, the node issues a new reset and returns, waiting for it to take place. Finally, if there is no primary block entry then this is the first execution of the protocol, so the node issues a reset and returns, waiting for it to take place.
@@ 5.4.1. Sync to Latest Checkpoint
5.4.2. Sync Past Checkpoint
@@ 5.4.3. Extend the log
5.5. Forensics
Aegis takes advantage of the consensus protocol’s forensics support (Sheng et al., 2021; Civit et al., 2021) to penalize Byzantine nodes if their number is greater than its threshold, and they successfully violate agreement. The forensics procedure encompasses both the contract and the peer-to-peer protocol, we describe it only here for pseudocode readability.
@@ 5.5. Forensics
6. Security
To prove Aegis’s security, we first show that the ancestors of a block generated by an active committee were also generated by active committees (§6.1). After that, we prove Aegis achieves agreement (§6.2), validity (§6.3) and progress (§6.4), and that it penalizes misbehaving nodes (§6.5). The security guarantees follow (§6.6). Appendix B summarizes the assumptions Aegis requires.
6.1. Active Committee Follows Active Committees
We call a committee active if its nodes’ collateral was locked recently enough, so it could not be withdrawn yet.
@@ 6.1. Active Committee Follows Active Committees
6.2. Agreement
For Agreement, we need two more lemmas. First, we ensure steps taken by the nodes are reflected in the primary chain. We show that if the nodes agree on the ledger up to some height, then a subsequent block will be accepted by the primary-chain contract.
@@ 6.2. Agreement
6.3. No-Batching Validity
To prove no-batching validity, we first note that the consensus protocol guarantees validity of an active committee decision. The proof is similar to that of Lemma 5.
@@ 6.3. No-Batching Validity
We are now ready to prove no-batching validity.
Proposition 8 (No-batching validity).
Aegis achieves no-batching validity.
@@
Assumption. For $k>0$, assume that the claim holds for ${k-1}$. That is, for all nodes $i,i^{\prime}$ (perhaps $i=i^{\prime}$), if node $i^{\prime}$ logs a block $b$ at height $k^{\prime}\leq k-1$ and node $i$ logs a block at height $k^{\prime}-1$ at time $t$ and all nodes are correct with input $v$ from time $t$ ($\forall j,t^{\prime\prime}>t:v^{j}_{t^{\prime\prime}}(k^{\prime})=v_{k^{\prime}}$) then node $i^{\prime}$ logs $v$ at height $k^{\prime}$${\mathcal{L}_{i^{\prime}}^{t_{1}}(k^{\prime})=v_{k^{\prime}}}$. Also assume there is no checkpoint for a block $b$ at a height $k^{\prime}\leq k-1$ with value different from $v_{k^{\prime}}$.
Step. Now we prove for $k$. Consider a node $i$ and time $t$ such that node $i$ logs a block $b$ in position $k$ at time $t$ and $\mathcal{L}_{i}^{t_{1}}(k)=b\neq v_{k}$.
@@
In both cases, the active committee nodes call the input function only after observing the ${k-1}$ block (lines 4 or 4). Thus, in both cases, all inputs of the consensus nodes are read from the input function $v(\cdot)$ after they logged block ${k-1}$. If the input function returns the same value $v_{k}$ for all correct nodes after the first decision on block ${k-1}$, this will be the input of all the active-committee nodes for block $k$. The consensus validity property (Lemma 7) guarantees that the consensus protocol outputs this value and this is their decision.
@@ 6.4. Progress
Consider the following worst case scenario. At time $t>t_{\textit{GST}}$ the Aegis checkpoint is later than $t-\Delta_{\textit{active}}+3\Delta_{\underline{\mathcal{L}}\textit{-write}}$, the nodes take consensus steps and decide within $\Delta_{\textit{consensus}}$ steps (Algorithm 5.3 line 4, Lemma 9). However, if the consensus runs for too long, and its result might not be checkpointed in time, it is abandoned (Algorithm 5.3 line 4). Then no progress can be made until $\Delta_{\textit{active}}$ after the last checkpoint/reset and a new committee is instated with a new reset (line 4). The reset is registered and nodes see it within $2\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps, and a new committee reaches consensus within $\Delta_{\textit{consensus}}$ (Lemma 9).
Other scenarios only speed up progress: If the checkpoint is earlier, the reset will be issued faster, and if the consensus following the checkpoint succeeds the block will be created faster.
Overall, the next block is created at most $\Delta_{\textit{stabilize}}=\Delta_{\textit{consensus}}+\Delta_{\textit{prop}}+\Delta_{\textit{active}}+2\Delta_{\underline{\mathcal{L}}\textit{-write}}$ steps after $t_{\textit{GST}}$.
@@
6.5. Slashing
Enforcing the desired behavior of correct nodes while they are active necessitates penalizing them if they successfully violate safety (Agreement or Validity). We show Aegis achieves this if the time it takes to unstake is long enough.
@@ 6.5. Slashing
Assume $\Delta_{\textit{active}}>3\Delta_{\underline{\mathcal{L}}\textit{-write}}$. If a node $i$ belongs to an active committee that violates the consensus safety, it will be penalized.
@@
6.6. Aegis Security
The conclusion follows directly from Propositions 68, and 10, since nodes are penalized (Proposition 11) for violating safety.
@@ 6.6. Aegis Security
Aegis ensures Agreement, No-Batching Validity, and Progress.
7. Practical Considerations and Broader Applicability
While Aegis can be deployed as-is on top of a primary blockchain, there are two practical aspects to consider. Aegis requires primary-chain transactions in intervals close to the time for unstaking $\Delta_{\textit{active}}$. This implies a tradeoff between short unstaking time and low primary-chain overhead.
In terms of performance, Aegis can change committee membership at a lower frequency than every block. This allows using contemporary BFT state machine replication algorithms (e.g., (Yin et al., 2019; Gelashvili et al., 2022; Babel et al., 2023)) that are more efficient than running a one-shot consensus instance per block.
Beyond stake-based systems, our tethering approach provides a principled solution for secure epoch changes in blockchains more broadly. Aegis can be adapted to support various committee transition criteria instead of staking and unstaking. For example, it could trigger based on a certain number of blocks (time-based epochs), based on consensus of a subcommittee, or by a central authority. To implement such extensions, the smart contract on the primary chain would need additional logic to verify the transition conditions before accepting resets. In addition, this would require tailoring the node protocol to match the transition logic in the smart contract in order to maintain the security guarantees of Aegis.
8. Conclusion
We present Aegis, a novel tethered blockchain protocol whose nodes can join and leave in a permissionless fashion by staking through a primary-chain smart contract. Aegis guarantees safety at all times and progress starting shortly after $t_{\textit{GST}}$ despite asynchronous communication among its nodes up to $t_{\textit{GST}}$. A core assumption allowing Aegis to succeed is that the nodes have synchronous access to the primary chain. Without this assumption, no protocol can achieve all desired properties: Agreement, No-Batching Validity and Progress.
The tools to deploy staking for Aegis are ready at hand, e.g., Avalanche subnets (Labs, 2025); direct implementation with a smart contract; and Ethereum’s EigenLayer (EigenLayer, 2023), which enables a market for stake on Ethereum that secures external tasks as well as Ethereum’s consensus. Thus, Aegis opens the door for a new generation of tethered blockchains supporting truly decentralized applications.
Appendix A Notation
Table 1. Notation
Symbol Meaning
$\alpha$ Minimum ratio of correct nodes in an active committee
$t_{\textit{GST}}$ Global Stabilization Time
$\Delta_{\textit{prop}}$ Network propagation bound after $t_{\textit{GST}}$
$H(B)=\hat{B}$ Hash of block $B$
$\underline{\mathcal{L}}$ Primary ledger
$\mathcal{L}$ Aegis ledger
$\mathcal{L}(k)$ Aegis block at height $k$
$\textit{parent}(b)$ Parent block of block $b$
$\mathcal{B}(\hat{b})$ Block $b$ if $b$ is in block set $B$, otherwise $\bot$
$\textit{ref}_{A\to P}(b)$ Reference from block $b$ to a primary block
$\textit{stakers}(B)$ Committee defined by primary block $B$
$t_{\textit{gen}}(b)$ Generation time of block $b$
Table 1 summarizes key notation used in this paper.
Appendix B Security Assumptions
To provide a comprehensive view of Aegis’s security model, we reiterate its security assumptions.
  1. (1)
    Primary Blockchain Properties (§3.1)
    1. (a)
      Primary blockchain security. The primary chain is implemented by a trusted party that maintains token balances and staked amounts and supports smart contracts.
    2. (b)
      Unpredictable block hashes. Nodes cannot guess future primary chain block hashes except with negligible probability.
  2. (2)
    Node Behavior and Stake Properties (§3.1)
    1. (a)
      Sufficient unstaking delay. Unstaking completes $\Delta_{\textit{active}}$ steps after the request transaction is placed.
    2. (b)
      Honest majority in committees. In all committees, the ratio of stake held by correct nodes exceeds $\alpha$.
    3. (c)
      Stake-based correctness incentive. Correct nodes follow the protocol while staked to avoid losing stake.
  3. (3)
    Network Assumptions (§3.1)
    1. (a)
      Synchronous primary chain access. Nodes can observe state and issue transactions within bounded time $\Delta_{\underline{\mathcal{L}}\textit{-write}}$.
    2. (b)
      Partially synchronous node communication. After $t_{\textit{GST}}$, message delivery between nodes is bounded by $\Delta_{\textit{prop}}$.
    3. (c)
      Block propagation persistence. Blocks are delivered to all nodes, even those staking later.
  4. (4)
    Consensus Algorithm Properties (§3.2)
    1. (a)
      Underlying consensus guarantees. If the ratio of correct nodes in a committee exceeds $\alpha$, the consensus protocol ensures agreement (no two correct nodes decide different values), validity (if all nodes are correct with same input, no other value is decided), and termination (after $t_{\textit{GST}}$, all correct nodes decide within $\Delta_{\textit{consensus}}$ time).
    2. (b)
      Consensus validation function. There exists a function consensusValidate that verifies block validity.
    3. (c)
      Forensics support. The consensus protocol allows identifying Byzantine nodes.
  5. (5)
    Timing Constraints (§3.2)
    1. (a)
      Minimum unstaking delay. Unstaking is slow enough compared to the time to issue a transaction to the primary chain: ${\Delta_{\textit{active}}>3\Delta_{\underline{\mathcal{L}}\textit{-write}}}$.